Calculate the beta for Stock B. Consider a portfolio that is one-third invested in Stock A and one-third invested in Stock B, as well as, one-third invested in a risk-free asset. Stock A has a beta of 1.54. The total portfolio parallels the risk to the market. Download the linked spreadsheet template and use it for your answer. Once complete, upload the template to this question. Don't forget to show your work!

Input area:
Weight of risk-free 33.33%
Weight of Stock A 33.33%
Weight of Stock B 33.33%
Beta of risk-free 0.00
Beta of Stock A         1.54
Beta of Portfolio         1.00
Output area:
Beta of Stock B
Answer #1

E18 fx А B с D E 1 2 Weight Beta Weight * Beta 3 0.33 0 0 Particualrs Risk Free rate Stock A Stock B 4 0.5082 5 0.33 1.54 0.3E18 А D F 1 2 Beta 0 3 4 Particualrs Risk Free rate Stock A Stock B Weight 0.33 0.33 0.33 Weight * Beta =C3*D3 =C4*D4 =E6-E4-